a car travelling at velosity of 200m/s for 2 sec in nesxt 5 secs it becomes the 220m/s then what is the velocity of car at the end of 7th second from start?

Grade:9

1 Answers

sonika p
6 years ago
for 1st  2 sec, 
v=u+at, v=220m/s, u=200m/s
220=200+a(2)
a=10m/s2.
for next 5sec
v=u+at, u=220m/s, a=10m/s2, t=5sec
v=220+(10*5)
v=270m/s
hence, velocity of car at the end of 7th second 270m/s.